hapter 2 Question 12 (of 12) sh $320,000; Determine the common stock for Bertinell Corp based on the following information: ca patents and copyrights $650000. accounts payable-$420.000, accounts recelvable-$119,000, tangible net fixed assets $4.800,000 inventory $285,000, notes payable-$170,000; accumulated retained earnings $1.305,000, long-term debt $1930000 O $2.349.000 O $2.689.000 O $2.097000 ○ S2601000 O $2.951.000 Ask me anything f4 fs 19 72345 67 8 9
Answer #1

Total assets=Cash+patents and copyrights+AR+Tangible net fixed assets+inventory

=(320,000+650,000+119000+4,800,000+285000)=$6,174,000

Total liabilities=AP+notes payable+long term debt

=(420,000+170,000+1930000)=$2,520,000

Total assets=Total liabilities+Total equity

6,174,000=2,520,000+(Common stock+1305000)

Common stock=(6174000-2520000-1305000)

which is equal to

=$2,349,000